Pete Pablo has $20000 to invest. He is very optimistic about the prospects of two companies, 919 Brands and Diaries.com. However, Pete has a very pessimistic view of one company, a financial institution known as Star Bank. The current market price of each share and Pete’s assessment of the expected returns for each share are below

EvSJkAAAAAElFTkSuQmCC

a) Pete decides to purchase 210 shares of 919 Brands and 180 shares of Diaries.com. What is the expected return on his portfolio? Can Pete construct this portfolio with the amount of money he has?
b) If Pete sells short 100 shares of Star Bank, how much additional money will he have to invest in the other two shares?
c) If Pete buys 210 shares of 919 Brands and 180 shares of Diaries.com, and he simultaneously sells short 100 shares of Star Bank, what are the resulting portfolio weights in each share? What is the expected return on the portfolio?
Answer #1
a. Total investment (inthis portfolio)based on his decision:
(210*60)+(180*80)=
27000
Expected return on the above portfolio=(Sum of Weights(value) to total*Respective exp.return) --of both stocks
ie((210*60)/27000*10%)+((180*80)/27000*14%)=
12.13%
NO
Reqd. 27000 > 20000 he has.
b. By selling short, 100 shares of Star Bank, he will have an additional money of
100*70= 7000
to invest in the other two stocks.
c. Expected return on the portfolio:
No.of shares Price/share Total value Wts. Of values to total value Exp.Ret. Wts. *exp.Ret.
210 60 12600 63.00% 10% 6.30%
180 80 14400 72.00% 14% 10.08%
100 0.00% -8% 0.00%
-100 70 -7000 -35.00% -8% 2.80%
20000 1 19.18%
Expected return= 19.18%